OpenStudy (anonymous):

1. A philanthropist gives $50,000 to students who have helped the community. The money is divided equally among these students. There is no limit to the number of students x. All students with good community projects will get a share of the money. In addition, these students will get $2,000 up front from a separate fund. What equation shows the amount of money each student receives? A.) Y = 50,000/x+2,000. B.) Y = 50,000/ x + 2,000. C.) Y = 50,000/x-2,000. D.) Y = 50,000/ x - 2,000. 2. A trip from Ohio to New York is 529 miles. What equation shows the time t it takes to go by car if

OpenStudy (anonymous):

the average speed during the trip? A.) t = 529/r B.) t = 529r C.) r = 529t D.) t = r/529

OpenStudy (anonymous):

Y=(50,000/x) + 2000...because 50,000 is reparted between x students and 2000 comes from a separate fund

OpenStudy (anonymous):

4. Describe how the graph of g(x)= 3/x - 4 + 5 is a translation of g(x)= 3/x . A.) It is a translation of g(x)= 3/x , 4 units right and 5 units down. B.) It is a translation of g(x)= 3/x, 4 units right and 5 units up. C.) It is a translation of g(x)= 3/x , 4 units left and 5 units up. D.) It is a translation of g(x)= 3/x , 4 units left and 5 units down.

OpenStudy (anonymous):

would that be A or B? @gerlpz22

OpenStudy (anonymous):

5. Identify the asymptotes of the graph. A.) x = –2, y = –1 B.) x = 2, y = 1 C.) x = 1, y = –2 D.) x = –1, y = 2

OpenStudy (anonymous):

I need help. :(

OpenStudy (anonymous):

a

OpenStudy (anonymous):

for question 5. is B

OpenStudy (anonymous):

in question 4...the -4 is part of the denominator?

OpenStudy (anonymous):

yes @gerlpz22

OpenStudy (anonymous):

what is #4? @gerlpz22